I am having a problem with an equation but cannot figure out how to find it. The problem is: 30^2 = ((s+7)(1.5))^2 + (2s)^2
The answer for "s" is 9. What am I doing wrong?
(30)(30) = ((1.5s)+(1.5)(7))^2 + (2s)(2s)
900 = (1.5s + 10.5)^2 + 4s^2
900 = ((1.5s)(1.5s)+ (10.5)(10.5)) + 4s^2
900 = 2.25s^2 + 110.25 + 4s^2
900 - 110.25 = 2.25s^2 +110.25-110.25 + 4s^2
789.75 = 6.25s^2
789.75/6.25 = 6.25s^2/6.25
126.36 = s^2
sqrt 126.36 = s ---sqrt 126.36 does not equal 9---

Answer by Alan3354(21583)

I am having a problem with an equation but cannot figure out how to find it. The problem is: 30^2 = ((s+7)(1.5))^2 + (2s)^2
The answer for "s" is 9. What am I doing wrong?
(30)(30) = ((1.5s)+(1.5)(7))^2 + (2s)(2s)
900 = (1.5s + 10.5)^2 + 4s^2
***************** You left out the s term in (1.5s + 10.5)^2
** (1.5s + 10.5)^2 = 2.25s^2 + 2*1.5*10.5s + 110.25
***********
900 = ((1.5s)(1.5s)+ (10.5)(10.5)) + 4s^2
900 = 2.25s^2 + 110.25 + 4s^2
900 - 110.25 = 2.25s^2 +110.25-110.25 + 4s^2
789.75 = 6.25s^2
789.75/6.25 = 6.25s^2/6.25
126.36 = s^2
sqrt 126.36 = s ---sqrt 126.36 does not equal 9---
---------------------------